OpenStudy (anonymous):

using the quadratic formula, solve the quadratic equation: 6x^2-7x-3=0

OpenStudy (anonymous):

I know how to fill it in... I need help getting the actual answers because they will only come out as weird decimals

OpenStudy (anonymous):

I am getting \[ x = \frac{1}{12} (7 \pm 11)\] so \[ x =- \frac{1}{3} \] and \[ x = \frac{3}{2} \] see step by step work in the attached file confirmed solution: http://www.wolframalpha.com/input/?i=6*x%5E2-7*x-3+%3D+0